The school purchased baseball equipment and uniforms for a total cost of $1,762.00. The equipment cost $598.00 and the uniforms were $24.25 each. How many uniforms the school purchase?
Answer by checkley71(8403)

1762-598=24.25X OR 1164=24.25X OR X=1164/24.25 OR X=48 UNIFORMS PURCHASED FOR 24.25 EACH